Give the percent yield when 14.08G of C02 are formed from the reaction of 2.000 moles of C8H 18 with 8.000 moles of 02.

2 C8H18 + 25 02 ——> 16 C02 +18 H2O
Answer #1

Balanced chemical reaction is:

2 C8H18 + 25 02 ——> 16 C02 +18 H2O

moles of O2 needed to react with 2 mol of C8H18 = (25 / 2) x 2 mol = 25 mol

Here; moles of O2 reacted is = 8.0 mol (Oxygen present in limited quantity)

So; the limiting reagent is O2

moles of CO2 produced from 8.0 mol of O2 = (16 / 25) x 8 mol = 5.12 mol

Theoretical weight of CO2 = (5.12 mol) x (44.01 g/mol) = 225.33 g

practical yield of CO2 = 14.08 g

Percent yield = (14.08 / 225.33) x 100 = 6.25 %